OpenStudy (anonymous):

Twelve cells, each of EMF 2 volts and internal resistance 0.5 Ω, are connected in series across an external resistance of 4.5 Ω. Find the current in the circuit.

OpenStudy (anonymous):

i=2/(4.5+.5) i=2/5 A